Science Report – February 17th

SOL 1

Crew 222 Science Report 17Feb2020
Author : Nina SEDBON
Title : Several experiments begin

In the Morning :

• LOAC and Weather Station: Put in place during the EVA. They were working just fine.
• For the LOAC the conditions were easy to obtain : stable floor, parallel to the floor for the wind captor.
• For the weather station : Girouette North orientated, three feet at a 90° angle from the ground for the whole to be very straight and vertical.

In the Afternoon:

• Aerospring : the conditions are very different in the GreenHab and in the Science Dome. We need more information to start building it because it will be very difficult to change its location once it is built. We only planted : Swiss chard, Radish, Spinach and Peas to transfer it in the vertical garden after it starts growing.
• MegaAres: Few tries in the Lower Deck to prepare the EVA for tomorrow.
• In the GreenHab : Start of the Spiruline experiment in the two aquariums.
• Backpack : Few operations in the RAM to connect the LCVG to the backpack.
• In the RAM : Fixing of the hygrometric captor for the Weather Station.
• Human Factors experiment called Foreign Languages at 16h45. It is organized with sessions of two people for 15 minutes each time. The first session is in Spanish. The two others are in English. It is a cooperation game with LEGO.
• Water Experiment: Few checks for the electrical setup. We hope to start the experiment tomorrow night. We will describe the set up once it is put in place.

Problems faced today concerning the environment of the experiments: The lack of screen for the water experiment was unexpected. A computer has to be
used to replace it.
The lack of « oscillator » for the MegaAres. It is thus complicated to test it before the EVA.

Astronomy Report – February 17th

Crew 222 Astronomy Report Feb 17 2020

Name: Eishi Kim
Crew: 222
Date: 2/17/2020

MDRS ROBOTIC OBSERVATORY
Robotic Telescope Requested (choose one) MDRS-14

Objects to be Imaged this Evening: as many as possible among NGC 772, NGC 5248, IC 342, NGC 2775, NGC 4449, NGC 4254, NGC 4038, NGC 4125, NGC 3338, NGC 2683, M 100. This is part of the weekly observations of select galaxies as per the submitted research proposal on supernovae.

Images submitted with this report: N/A

Problems Encountered: N/A

MUSK OBSERVATORY

Solar Features Observed: the observatory was first used today. Attempts have been made to visually observe solar features according to the quick guide, but no special phenomenon was to be noted. This should not be an issue of the telescope but most certainly due to the weak solar activity.

Images submitted with this report: N/A

Problems Encountered: N/A

Journalist Report – February 17th

Crew 222 Journalist Report 17Feb2020
Author: Marie Bochard, Crew Journalist
Sol 1
Title: Problem-solving

Here on Mars, every small problem can take incredible proportions. Consider this: what if one life-support system fails? The entire mission is at stake, so everyone has to chip in and help repair the problem. Thankfully, this rarely happens, but minor problems still occur.

This morning was our first EVA, and there were two important experiments to get ready. The LOAC is used to measure air particles, which will be necessary to analyze the atmosphere on Mars, and the weather station can help predict dust storms that will be frequent on the red planet. Both experiments were assembled before leaving the Hab, because when we are on EVA we wear big gloves that really impair on our ability to manipulate tools. Once the setup was done, the four crew members leaving for EVA got ready in their spacesuits. These backpacks are designed specifically to enable humans to walk on Martian soil even though the environment is hostile: there is no oxygen on the surface of Mars, so us humans cannot breathe. This is an example of a life-support system that can need repair, since we cannot leave the station without it!

After a (very) spicy lunch, it was time for everyone to focus on their experiments. Valentin, the engineer, was tasked with finding a way to power a data collector for the weather station. This proved to be more difficult than we thought: after trying to solder batteries together to provide enough power to the data collector, which took us a few hours, we settled on using a powerbank. Problem-solving isn’t always about trying to make an idea work, it’s also about finding new ideas, and today was a proof for that.

Enough problems for now, today was also filled with crew moments. Meals are always a great way of communicating between us, since we all eat together. There moments are crucial for crew cohesion, because we are going to be living together for two weeks, so being honest and understanding everyone’s concern is fundamental. Cohesion is about sharing our favorite moments from every day, but also about making sure everyone feels at home and is their best self.

As the day winds down, we are all happy to smell the delicious blueberry tart that Bleuenn baked this afternoon. This has been an incredible first sim day, and we are looking forward to tomorrow!

Operations Report – February 17th

Crew 222 Operations Report 02/17/2020

SOL: 1

Name of person filing report: Valentin PRUDHOMME

Non-nominal systems: N/A

Notes on non-nominal systems: N/A

Generator: yes
Hours run: 11 hours
From what time last night: 7pm
To what time this morning: 6am
List any additional daytime hours when the generator was run: N/A

Solar—SOC 93% (Before generator is run at night)

Diesel Reading – 78%

Station Propane Reading – 41%

Ethanol Free Gasoline – 0.5 gallons

Water (loft tank): 46 gallons

Water Meter: 149275.24 units

Water (static tank) – 441 gallons

Static to Loft Pump used – yes

Water in GreenHab – 138 gallons

Water in ScienceDome: 0 gallons

Toilet tank emptied: no

Deimos rover used: yes
Hours: 143.7
Beginning charge: 100
Ending charge: 100
Currently charging: yes

Sojourner rover used: ASSIGNED TO DIRECTOR
Hours: N/A
Beginning charge: N/A
Ending charge: N/A
Currently charging: N/A

Spirit rover used: yes
Hours: 131.7
Beginning charge: 100
Ending charge:96
Currently charging: yes

Opportunity rover used: no
Hours: 69.9
Beginning charge: N/A
Ending charge: N/A
Currently charging: yes

Curiosity rover used: yes
Hours: 137.9
Beginning charge: 100
Ending charge: 91
Currently charging: yes

Notes on rovers: N/A

ATV’s Used: N/A
Reason for use: N/A
Oil Added? no
ATV Fuel Used: N/A
Hours the ATVs were Used today: 0
Notes on ATVs: N/A

HabCar used and why, where? N/A

CrewCar used and why, where? N/A

General notes and comments: N/A

Summary of internet: used for the com

Summary of suits and radios: Suits 7, 8, 9, 10 used and radio tested for the first EVA. All ok.

Summary of Hab operations: N/A

Summary of GreenHab operations: N/A

Summary of ScienceDome operations: N/A

Summary of RAM operations: Repaired the hygrometric captor of the weather station (finished). Working on the experiment of the backpack (not fully done).

Summary of any observatory issues: N/A

Summary of health and safety issues: N/A

Questions, concerns and requests to Mission Support: N/A

Greenhab Report – February 17th

Crew 222 GreenHab Report – 16/02/2020

GreenHab Officer : Bleuenn ROIRON

Environmental control : Ambient with door open

Average temperatures: 22°C

Hours of supplemental light: NA

Daily water usage for crops: 11,6 Gallons

Daily water usage for research and/or other purposes: 5 gallons

Water in blue Tank: 138,4 gallons

Time(s) of watering crops: 2

Changes to crops: NA

Narrative: Today, I watered plants before my first EVA and then asked to our journalist to take care of the greenhab,
This afternoon I went there to start my spiruline experiment, first i prepared the aquariums with the right components
and then I looked after plants. Nina, our crew scientist planted her plants(swiss chard, raddish, spinach, peas) that will be next planted in our culture
tower and we planted our first flower(zinnia) to add some colors in our simulation! Finally, tonight after having watered plants again I did injected spirulina in
the aquariums and my first experiment have started. I also harvested some aromatic herbs to cook tonight and make
our meals be more tasty!

Harvest: 5g lemon balm/ 4g basil/ 5g chives

Support/supplies needed: don’t have the real amount of water in the blue tank since i don’t know the real amount
that were in this tank at the beginning.

EVA Report – February 17th

Crew 222 EVA Report 17-02-2020
EVA # 1
Author: Benjamin Auzou
Purpose of EVA: Install weather station and aerosol counter, then find a place the installation of an other experiment, The electric field measure. Start time: 9:37am
End time: 11:33am
Narrative:
First of all this EVA was followed by a journalist, who filmed the activities on the field.
After the first steps on Mars for three of us, we checked all rovers status and the static tank level. Then we moved to the place we needed to set the experiment, which needs to be a plane area.
The installation was successful for both parts of the experiment. At 11:10am, the installation was finished and then we had some minutes to study the area to find a location for MegaAres, the experiment that we will deploy tomorrow.

Destination: plane area approximately 500m away from the Hab, along Cow Dung Road
Coordinates (use UTM NAD27 CONUS): 12S 0518483 4250677
Participants: Benjamin Auzou (CO), Bleuenn Roiron (GHO), Valentin Prudhomme (ENG), Nina Sedbon (SC)
Road(s) and routes per MDRS Map: Cow Dung Road
Mode of travel: Driving (needed for material transport)

Sol Summary – February 17th

Crew 222 Sol Summary Report 17-02-2020
Sol:1
Summary Title: And so the science begin
Author’s name: Benjamin Auzou
Mission Status: Crew settled and Beginning of the first experiments Sol Activity Summary: In the morning, the crew woke up with a matinal workout session. Then four crewmember went on the first EVA of the mission to install a field experiment.
The afternoon were dedicated to the launch of scientific experiments : Spirulina in the Green Hab, Human factors, preparation of the water consumption monitoring.
Between 8:30am and 5pm, the crew were followed by a journalist for a French TV channel. The journalist were respectful of the place and of our work.
Look Ahead Plan:
Tomorrow, we will go on a similar EVA to this morning, in order to set up an electric field mesure, MegaAres. Also we will continue the activities launched today.
Anomalies in work: none
Weather: cloudy in general, clearer in the afternoon
Crew Physical Status: all members safe and sound
EVA: LOAC and Weather station installation success
Reports to be filed: Journalist report, Mission Plan, Crew bios, GreenHab Report, Astronomy Report, EVA Report, EVA Request, Science report, HSO Pre Mission Checklist
Support Requested: Food Inventory to be reviewed by Shannon